Find coordinates of the foot of perpendicular from the point (3,-4) to line 4x – 15y + 17 = 0.

Grade:

1 Answers

SHAIK AASIF AHAMED
9 years ago
Hello student,
given equation is 4x-15y+17=0.......(1)
from the given equation the slope is 4/15
let pt be (x,y) and pt given (3,-4)
foot of perpendicular slope=(y-3)/(x+4)
as their slopes are perpendicular
(y-3)/(x+4)*4/15=-1...........(2)
by solving 1 and 2 we get foot of perpendicular as (-3.27,0.26)
Thanks and regards
Shaik Aasif
askIITians faculty